Glycerine, which is used for example in some cough syrups, toothpastes, skin care products, amongst others, has the formula C3H5(OH)3. Total mass of one mole is 92.094 g. If carbon, and oxygen have a molar mass of 12.01 g/mol and 16.00 g/mol, respectively, then what is the molar mass of hydrogen? Keep answer to the tenth. Include unit symbol.
1
Expert's answer
2020-12-14T14:35:25-0500

Mr(H)=(Mr(C3H5(OH)3) - 3*Mr(C) - 3*Mr(O))/8=(92.094 - 36.03 - 48.00)/8 = 1.008 g/mol